GNU Grep: Print lines that match patterns

version 3.11, 10 May 2023

Alain Magloire et al.

This manual is for grep, a pattern matching engine. Copyright \begin{matrix}c \\ ◯ \end{matrix} 1999–2002, 2005, 2008–2023 Free Software Foundation, Inc.

Permission is granted to copy, distribute and/or modify this document under the terms of the GNU Free Documentation License, Version 1.3 or any later version published by the Free Software Foundation; with no Invariant Sections, with no Front-Cover Texts, and with no Back-Cover Texts. A copy of the license is included in the section entitled “GNU Free Documentation License”.

Table of Contents

1 Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Invoking grep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

2.1 Command-line Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.1.1 Generic Program Information . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.1.2 Matching Control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.1.3 General Output Control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.1.4 Output Line Prefix Control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.1.5 Context Line Control . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 2.1.6 File and Directory Selection. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 2.1.7 Other Options . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9

2.2 Environment Variables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 2.3 Exit Status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 2.4 grep Programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

3 Regular Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

3.1 Fundamental Structure . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 3.2 Character Classes and Bracket Expressions . . . . . . . . . . . . . . . . . . . . 15 3.3 Special Backslash Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 3.4 Anchoring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.5 Back-references and Subexpressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 3.6 Basic vs Extended Regular Expressions . . . . . . . . . . . . . . . . . . . . . . . . 17 3.7 Problematic Regular Expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 3.8 Character Encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 3.9 Matching Non-ASCII and Non-printable Characters . . . . . . . . . . . . 20

4 Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

5 Performance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

6 Reporting bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

6.1 Known Bugs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27

7 Copying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

7.1 GNU Free Documentation License . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28

Index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 37

1 Introduction

Given one or more patterns, grep searches input files for matches to the patterns. When it finds a match in a line, it copies the line to standard output (by default), or produces whatever other sort of output you have requested with options.

Though grep expects to do the matching on text, it has no limits on input line length other than available memory, and it can match arbitrary characters within a line. If the final byte of an input file is not a newline, grep silently supplies one. Since newline is also a separator for the list of patterns, there is no way to match newline characters in a text.

2 Invoking grep

The general synopsis of the grep command line is

     grep [option...] [patterns] [file...]

There can be zero or more option arguments, and zero or more file arguments. The patterns argument contains one or more patterns separated by newlines, and is omitted when patterns are given via the ‘-e patterns’ or ‘-f file’ options. Typically patterns should be quoted when grep is used in a shell command.

2.1 Command-line Options

grep comes with a rich set of options: some from POSIX and some being GNU extensions. Long option names are always a GNU extension, even for options that are from POSIX specifications. Options that are specified by POSIX, under their short names, are explicitly marked as such to facilitate POSIX-portable programming. A few option names are provided for compatibility with older or more exotic implementations.

Several additional options control which variant of the grep matching engine is used. See Section 2.4 [grep Programs], page 12.

2.1.1 Generic Program Information

--help Print a usage message briefly summarizing the command-line options and the bug-reporting address, then exit.

-V
--version

Print the version number of grep to the standard output stream. This version number should be included in all bug reports.

2.1.2 Matching Control

-e patterns
--regexp=patterns

Use patterns as one or more patterns; newlines within patterns separate each pattern from the next. If this option is used multiple times or is combined with the -f (--file) option, search for all patterns given. Typically patterns should be quoted when grep is used in a shell command. (-e is specified by POSIX.)

-f file
--file=file

Obtain patterns from file, one per line. If this option is used multiple times or is combined with the -e (--regexp) option, search for all patterns given. When file is ‘-’, read patterns from standard input. The empty file contains zero patterns, and therefore matches nothing. (-f is specified by POSIX.)

-i
-y
--ignore-case

Ignore case distinctions in patterns and input data, so that characters that differ only in case match each other. Although this is straightforward when letters differ in case only via lowercase-uppercase pairs, the behavior is unspecified in other situations. For example, uppercase “S” has an unusual lowercase counterpart “” (Unicode character U+017F, LATIN SMALL LETTER LONG S) in many locales, and it is unspecified whether this unusual character matches “S” or “s” even though uppercasing it yields “S”. Another example: the lowercase German letter “ß” (U+00DF, LATIN SMALL LETTER SHARP S) is normally capitalized as the two-character string “SS” but it does not match “SS”, and it might not match the uppercase letter (U+1E9E, LATIN CAPITAL LETTER SHARP S) even though lowercasing the latter yields the former.

-y is an obsolete synonym that is provided for compatibility. (-i is specified by POSIX.)

--no-ignore-case

Do not ignore case distinctions in patterns and input data. This is the default. This option is useful for passing to shell scripts that already use -i, in order to cancel its effects because the two options override each other.

-v
--invert-match

Invert the sense of matching, to select non-matching lines. (-v is specified by POSIX.)

-w
--word-regexp

Select only those lines containing matches that form whole words. The test is that the matching substring must either be at the beginning of the line, or preceded by a non-word constituent character. Similarly, it must be either at the end of the line or followed by a non-word constituent character. Word constituent characters are letters, digits, and the underscore. This option has no effect if -x is also specified.

Because the -w option can match a substring that does not begin and end with word constituents, it differs from surrounding a regular expression with ‘\<’ and ‘\>’. For example, although ‘grep -w @’ matches a line containing only ‘@’, ‘grep '\<@\>'’ cannot match any line because ‘@’ is not a word constituent. See Section 3.3 [Special Backslash Expressions], page 16.

-x
--line-regexp

Select only those matches that exactly match the whole line. For regular expression patterns, this is like parenthesizing each pattern and then surrounding it with ‘^’ and ‘$’. (-x is specified by POSIX.)

2.1.3 General Output Control

-c

--count Suppress normal output; instead print a count of matching lines for each input file. With the -v (--invert-match) option, count non-matching lines. (-c is specified by POSIX.)

--color[=WHEN]
--colour[=WHEN]

Surround matched non-empty strings, matching lines, context lines, file names, line numbers, byte offsets, and separators (for fields and groups of context lines) with escape sequences to display them in color on the terminal. The colors are defined by the environment variable GREP_COLORS and default

to ‘ms=01;31:mc=01;31:sl=:cx=:fn=35:ln=32:bn=32:se=36’ for bold red

matched text, magenta file names, green line numbers, green byte offsets, cyan separators, and default terminal colors otherwise. See Section 2.2 [Environment Variables], page 9. WHEN is ‘always’ to use colors, ‘never’ to not use colors, or ‘auto’ to use colors if standard output is associated with a terminal device and the TERM environment variable’s value suggests that the terminal supports colors. Plain --color is treated like --color=auto; if no --color option is given, the default

is --color=never.
-L
--files-without-match

Suppress normal output; instead print the name of each input file from which no output would normally have been printed.

-l
--files-with-matches

Suppress normal output; instead print the name of each input file from which output would normally have been printed. Scanning each input file stops upon first match. (-l is specified by POSIX.)

-m num
--max-count=num

Stop after the first num selected lines. If num is zero, grep stops right away without reading input. A num of −1 is treated as infinity and grep does not stop; this is the default. If the input is standard input from a regular file, and num selected lines are output, grep ensures that the standard input is positioned just after the last selected line before exiting, regardless of the presence of trailing context lines. This enables a calling process to resume a search. For example, the following shell script makes use of it:

     while grep -m 1 'PATTERN'
     do
       echo xxxx
     done < FILE

But the following probably will not work because a pipe is not a regular file:

     # This probably will not work.
     cat FILE |
     while grep -m 1 'PATTERN'
     do
       echo xxxx
     done

When grep stops after num selected lines, it outputs any trailing context lines. When the -c or --count option is also used, grep does not output a count greater than num. When the -v or --invert-match option is also used, grep stops after outputting num non-matching lines.

-o
--only-matching

Print only the matched non-empty parts of matching lines, with each such part on a separate output line. Output lines use the same delimiters as input, and delimiters are null bytes if -z (--null-data) is also used (see Section 2.1.7 [Other Options], page 9).

-q
--quiet

--silent Quiet; do not write anything to standard output. Exit immediately with zero status if any match is found, even if an error was detected. Also see the -s or --no-messages option. Portability note: Solaris 10 grep lacks -q; portable shell scripts typically can redirect standard output to /dev/null instead of using -q. (-q is specified by POSIX.)

-s
--no-messages

Suppress error messages about nonexistent or unreadable files. (-s is specified by POSIX.)

2.1.4 Output Line Prefix Control

When several prefix fields are to be output, the order is always file name, line number, and byte offset, regardless of the order in which these options were specified.

-b
--byte-offset

Print the 0-based byte offset within the input file before each line of output. If -o (--only-matching) is specified, print the offset of the matching part itself.

-H
--with-filename

Print the file name for each match. This is the default when there is more than one file to search.

-h
--no-filename

Suppress the prefixing of file names on output. This is the default when there is only one file (or only standard input) to search.

--label=LABEL

Display input actually coming from standard input as input coming from file LABEL. This can be useful for commands that transform a file’s contents before searching; e.g.:

     gzip -cd foo.gz | grep --label=foo -H 'some pattern'

-n
--line-number

Prefix each line of output with the 1-based line number within its input file. (-n is specified by POSIX.)

-T
--initial-tab

Make sure that the first character of actual line content lies on a tab stop, so that the alignment of tabs looks normal. This is useful with options that prefix their output to the actual content: -H, -n, and -b. This may also prepend spaces to output line numbers and byte offsets so that lines from a single file all start at the same column.

-Z

--null Output a zero byte (the ASCII NUL character) instead of the character that normally follows a file name. For example, ‘grep -lZ’ outputs a zero byte after each file name instead of the usual newline. This option makes the output unambiguous, even in the presence of file names containing unusual characters like newlines. This option can be used with commands like ‘find -print0’, ‘perl -0’, ‘sort -z’, and ‘xargs -0’ to process arbitrary file names, even those that contain newline characters.

2.1.5 Context Line Control

Context lines are non-matching lines that are near a matching line. They are output only if one of the following options are used. Regardless of how these options are set, grep never outputs any given line more than once. If the -o (--only-matching) option is specified, these options have no effect and a warning is given upon their use.

-A num
--after-context=num

Print num lines of trailing context after matching lines.

-B num
--before-context=num

Print num lines of leading context before matching lines.

-C num
-num
--context=num

Print num lines of leading and trailing output context.

--group-separator=string

When -A, -B or -C are in use, print string instead of -- between groups of lines.

--no-group-separator

When -A, -B or -C are in use, do not print a separator between groups of lines.

Here are some points about how grep chooses the separator to print between prefix fields and line content:

• Matching lines normally use ‘:’ as a separator between prefix fields and actual line content.

• Context (i.e., non-matching) lines use ‘-’ instead.

• When context is not specified, matching lines are simply output one right after another.

• When context is specified, lines that are adjacent in the input form a group and are output one right after another, while by default a separator appears between nonadjacent groups.

• The default separator is a ‘--’ line; its presence and appearance can be changed with the options above.

• Each group may contain several matching lines when they are close enough to each other that two adjacent groups connect and can merge into a single contiguous one.

2.1.6 File and Directory Selection

-a

--text Process a binary file as if it were text; this is equivalent to the

‘--binary-files=text’ option.
--binary-files=type

If a file’s data or metadata indicate that the file contains binary data, assume that the file is of type type. Non-text bytes indicate binary data; these are either output bytes that are improperly encoded for the current locale (see Section 2.2 [Environment Variables], page 9), or null input bytes when the -z (--null-data) option is not given (see Section 2.1.7 [Other Options], page 9).

By default, type is ‘binary’, and grep suppresses output after null input binary data is discovered, and suppresses output lines that contain improperly encoded data. When some output is suppressed, grep follows any output with a message to standard error saying that a binary file matches.

If type is ‘without-match’, when grep discovers null input binary data it assumes that the rest of the file does not match; this is equivalent to the -I option.

If type is ‘text’, grep processes binary data as if it were text; this is equivalent to the -a option.

When type is ‘binary’, grep may treat non-text bytes as line terminators even without the -z (--null-data) option. This means choosing ‘binary’ versus ‘text’ can affect whether a pattern matches a file. For example, when type is ‘binary’ the pattern ‘q$’ might match ‘q’ immediately followed by a null byte, even though this is not matched when type is ‘text’. Conversely, when type is ‘binary’ the pattern ‘.’ (period) might not match a null byte.

Warning: The -a (--binary-files=text) option might output binary garbage, which can have nasty side effects if the output is a terminal and if the terminal driver interprets some of it as commands. On the other hand, when reading files whose text encodings are unknown, it can be helpful to use -a or to set ‘LC_ALL='C'’ in the environment, in order to find more matches even if the matches are unsafe for direct display.

-D action
--devices=action

If an input file is a device, FIFO, or socket, use action to process it. If action is ‘read’, all devices are read just as if they were ordinary files. If action is ‘skip’, devices, FIFOs, and sockets are silently skipped. By default, devices are read if they are on the command line or if the -R (--dereference-recursive) option is used, and are skipped if they are encountered recursively and the -r (--recursive) option is used. This option has no effect on a file that is read via standard input.

-d action
--directories=action

If an input file is a directory, use action to process it. By default, action is ‘read’, which means that directories are read just as if they were ordinary files (some operating systems and file systems disallow this, and will cause grep to print error messages for every directory or silently skip them). If action is ‘skip’, directories are silently skipped. If action is ‘recurse’, grep reads all files under each directory, recursively, following command-line symbolic links and skipping other symlinks; this is equivalent to the -r option.

--exclude=glob

Skip any command-line file with a name suffix that matches the pattern glob, using wildcard matching; a name suffix is either the whole name, or a trailing part that starts with a non-slash character immediately after a slash (‘/’) in the name. When searching recursively, skip any subfile whose base name matches glob; the base name is the part after the last slash. A pattern can use ‘*’, ‘?’, and ‘[’...‘]’ as wildcards, and \ to quote a wildcard or backslash character literally.

--exclude-from=file

Skip files whose name matches any of the patterns read from file (using wildcard matching as described under --exclude).

--exclude-dir=glob

Skip any command-line directory with a name suffix that matches the pattern glob. When searching recursively, skip any subdirectory whose base name matches glob. Ignore any redundant trailing slashes in glob.

-I Process a binary file as if it did not contain matching data; this is equivalent

to the ‘--binary-files=without-match’ option.
--include=glob

Search only files whose name matches glob, using wildcard matching as described under --exclude. If contradictory --include and --exclude options are given, the last matching one wins. If no --include or --exclude options match, a file is included unless the first such option is --include.

-r
--recursive

For each directory operand, read and process all files in that directory, recursively. Follow symbolic links on the command line, but skip symlinks that are encountered recursively. Note that if no file operand is given, grep searches the working directory. This is the same as the ‘--directories=recurse’ option.

-R
--dereference-recursive

For each directory operand, read and process all files in that directory, recursively, following all symbolic links.

2.1.7 Other Options

-- Delimit the option list. Later arguments, if any, are treated as operands even if they begin with ‘-’. For example, ‘grep PAT -- -file1 file2’ searches for the pattern PAT in the files named -file1 and file2.

--line-buffered

Use line buffering for standard output, regardless of output device. By default, standard output is line buffered for interactive devices, and is fully buffered otherwise. With full buffering, the output buffer is flushed when full; with line buffering, the buffer is also flushed after every output line. The buffer size is system dependent.

-U

--binary On platforms that distinguish between text and binary I/O, use the latter when reading and writing files other than the user’s terminal, so that all input bytes are read and written as-is. This overrides the default behavior where grep follows the operating system’s advice whether to use text or binary I/O. On MS-Windows when grep uses text I/O it reads a carriage return–newline pair as a newline and a Control-Z as end-of-file, and it writes a newline as a carriage return–newline pair. When using text I/O --byte-offset (-b) counts and --binary-files heuristics apply to input data after text-I/O processing. Also, the --binary-files heuristics need not agree with the --binary option; that is, they may treat the data as text even if --binary is given, or vice versa. See Section 2.1.6 [File and Directory Selection], page 7. This option has no effect on GNU and other POSIX-compatible platforms, which do not distinguish text from binary I/O.

-z
--null-data

Treat input and output data as sequences of lines, each terminated by a zero byte (the ASCII NUL character) instead of a newline. Like the -Z or --null option, this option can be used with commands like ‘sort -z’ to process arbitrary file names.

2.2 Environment Variables

The behavior of grep is affected by several environment variables, the most important of which control the locale, which specifies how grep interprets characters in its patterns and data.

The locale for category LC_foo is specified by examining the three environment variables LC_ALL, LC_foo, and LANG, in that order. The first of these variables that is set specifies the locale. For example, if LC_ALL is not set, but LC_COLLATE is set to ‘pt_BR.UTF-8’, then a Brazilian Portuguese locale is used for the LC_COLLATE category. As a special case for LC_MESSAGES only, the environment variable LANGUAGE can contain a colon-separated list of languages that overrides the three environment variables that ordinarily specify the LC_MESSAGES category. The ‘C’ locale is used if none of these environment variables are set, if the locale catalog is not installed, or if grep was not compiled with national language support (NLS). The shell command locale -a lists locales that are currently available.

The following environment variables affect the behavior of grep.

GREP_COLOR

This obsolescent variable interacts with GREP_COLORS confusingly, and grep warns if it is set and is not overridden by GREP_COLORS. Instead of

‘GREP_COLOR='color'’, you can use ‘GREP_COLORS='mt=color'’.
GREP_COLORS

This variable controls how the --color option highlights output. Its value is a colon-separated list of terminfo capabilities that defaults to

‘ms=01;31:mc=01;31:sl=:cx=:fn=35:ln=32:bn=32:se=36’ with the ‘rv’ and

‘ne’ boolean capabilities omitted (i.e., false). The two-letter capability names refer to terminal “capabilities,” the ability of a terminal to highlight text, or change its color, and so on. These capabilities are stored in an online database and accessed by the terminfo library. Non-empty capability values control highlighting using Select Graphic Rendition (SGR) commands interpreted by the terminal or terminal emulator. (See the section in the documentation of your text terminal for permitted values and their meanings as character attributes.) These substring values are integers in decimal representation and can be concatenated with semicolons. grep takes care of assembling the result into a complete SGR sequence (‘\33[’...‘m’). Common values to concatenate include ‘1’ for bold, ‘4’ for underline, ‘5’ for blink, ‘7’ for inverse, ‘39’ for default foreground color, ‘30’ to ‘37’ for foreground colors, ‘90’ to ‘97’ for 16-color mode foreground colors, ‘38;5;0’ to ‘38;5;255’ for 88-color and 256-color modes foreground colors, ‘49’ for default background color, ‘40’ to ‘47’ for background colors, ‘100’ to ‘107’ for 16-color mode background colors, and ‘48;5;0’ to ‘48;5;255’ for 88-color and 256-color modes background colors. Supported capabilities are as follows.

sl= SGR substring for whole selected lines (i.e., matching lines when the -v command-line option is omitted, or non-matching lines when -v is specified). If however the boolean ‘rv’ capability and the -v command-line option are both specified, it applies to context matching lines instead. The default is empty (i.e., the terminal’s default color pair).

cx= SGR substring for whole context lines (i.e., non-matching lines when the -v command-line option is omitted, or matching lines when -v is specified). If however the boolean ‘rv’ capability and the -v command-line option are both specified, it applies to selected non-matching lines instead. The default is empty (i.e., the terminal’s default color pair).

rv Boolean value that reverses (swaps) the meanings of the ‘sl=’ and ‘cx=’ capabilities when the -v command-line option is specified. The default is false (i.e., the capability is omitted).

mt=01;31 SGR substring for matching non-empty text in any matching line (i.e., a selected line when the -v command-line option is omitted, or a context line when -v is specified). Setting this is equivalent to setting both ‘ms=’ and ‘mc=’ at once to the same value. The default is a bold red text foreground over the current line background.

ms=01;31 SGR substring for matching non-empty text in a selected line. (This is used only when the -v command-line option is omitted.) The effect of the ‘sl=’ (or ‘cx=’ if ‘rv’) capability remains active when this takes effect. The default is a bold red text foreground over the current line background.

mc=01;31 SGR substring for matching non-empty text in a context line. (This is used only when the -v command-line option is specified.) The effect of the ‘cx=’ (or ‘sl=’ if ‘rv’) capability remains active when this takes effect. The default is a bold red text foreground over the current line background.

fn=35 SGR substring for file names prefixing any content line. The default is a magenta text foreground over the terminal’s default background.

ln=32 SGR substring for line numbers prefixing any content line. The default is a green text foreground over the terminal’s default background.

bn=32 SGR substring for byte offsets prefixing any content line. The default is a green text foreground over the terminal’s default background.

se=36 SGR substring for separators that are inserted between selected line fields (‘:’), between context line fields (‘-’), and between groups of adjacent lines when nonzero context is specified (‘--’). The default is a cyan text foreground over the terminal’s default background.

ne Boolean value that prevents clearing to the end of line using Erase in Line (EL) to Right (‘\33[K’) each time a colorized item ends. This is needed on terminals on which EL is not supported. It is otherwise useful on terminals for which the back_color_erase (bce) boolean terminfo capability does not apply, when the chosen highlight colors do not affect the background, or when EL is too slow or causes too much flicker. The default is false (i.e., the capability is omitted).

Note that boolean capabilities have no ‘=’... part. They are omitted (i.e., false) by default and become true when specified.

LC_ALL
LC_COLLATE

LANG These variables specify the locale for the LC_COLLATE category, which might affect how range expressions like ‘a-z’ are interpreted.

LC_ALL
LC_CTYPE

LANG These variables specify the locale for the LC_CTYPE category, which determines the type of characters, e.g., which characters are whitespace. This category also determines the character encoding. See Section 3.8 [Character Encoding], page 19.

LANGUAGE
LC_ALL
LC_MESSAGES

LANG These variables specify the locale for the LC_MESSAGES category, which determines the language that grep uses for messages. The default ‘C’ locale uses American English messages.

POSIXLY_CORRECT

If set, grep behaves as POSIX requires; otherwise, grep behaves more like other GNU programs. POSIX requires that options that follow file names must be treated as file names; by default, such options are permuted to the front of the operand list and are treated as options.

TERM This variable specifies the output terminal type, which can affect what the --color option does. See Section 2.1.3 [General Output Control], page 3.

The GREP_OPTIONS environment variable of grep 2.20 and earlier is no longer supported, as it caused problems when writing portable scripts. To make arbitrary changes to how grep works, you can use an alias or script instead. For example, if grep is in the directory ‘/usr/bin’ you can prepend $HOME/bin to your PATH and create an executable script $HOME/bin/grep containing the following:

     #! /bin/sh
     export PATH=/usr/bin
     exec grep --color=auto --devices=skip "$@"

2.3 Exit Status

Normally the exit status is 0 if a line is selected, 1 if no lines were selected, and 2 if an error occurred. However, if the -q or --quiet or --silent option is used and a line is selected, the exit status is 0 even if an error occurred. Other grep implementations may exit with status greater than 2 on error.

2.4 grep Programs

grep searches the named input files for lines containing a match to the given patterns. By default, grep prints the matching lines. A file named - stands for standard input. If no input is specified, grep searches the working directory . if given a command-line option specifying recursion; otherwise, grep searches standard input. There are four major variants of grep, controlled by the following options.

-G
--basic-regexp

Interpret patterns as basic regular expressions (BREs). This is the default.

-E
--extended-regexp

Interpret patterns as extended regular expressions (EREs). (-E is specified by POSIX.)

-F
--fixed-strings

Interpret patterns as fixed strings, not regular expressions. (-F is specified by POSIX.)

-P
--perl-regexp

Interpret patterns as Perl-compatible regular expressions (PCREs). PCRE support is here to stay, but consider this option experimental when combined with the -z (--null-data) option, and note that ‘grep -P’ may warn of unimplemented features. See Section 2.1.7 [Other Options], page 9. For documentation, refer to https://www.pcre.org/, with these caveats:

• ‘\d’ matches only the ten ASCII digits (and ‘\D’ matches the complement), regardless of locale. Use ‘\p{Nd}’ to also match non-ASCII digits. (The behavior of ‘\d’ and ‘\D’ is unspecified after in-regexp directives like ‘(?aD)’.)

• Although PCRE tracks the syntax and semantics of Perl’s regular expressions, the match is not always exact. For example, Perl evolves and a Perl installation may predate or postdate the PCRE2 installation on the same host, or their Unicode versions may differ, or Perl and PCRE2 may disagree about an obscure construct.

• By default, grep applies each regexp to a line at a time, so the ‘(?s)’ directive (making ‘.’ match line breaks) is generally ineffective. However, with -z (--null-data) it can work:

                   $ printf 'a\nb\n' |grep -zP '(?s)a.b'
                   a
                   b

But beware: with the -z (--null-data) and a file containing no NUL byte, grep must read the entire file into memory before processing any of it. Thus, it will exhaust memory and fail for some large files.

3 Regular Expressions

A regular expression is a pattern that describes a set of strings. Regular expressions are constructed analogously to arithmetic expressions, by using various operators to combine smaller expressions. grep understands three different versions of regular expression syntax: basic (BRE), extended (ERE), and Perl-compatible (PCRE). In GNU grep, basic and extended regular expressions are merely different notations for the same pattern-matching functionality. In other implementations, basic regular expressions are ordinarily less powerful than extended, though occasionally it is the other way around. The following description applies to extended regular expressions; differences for basic regular expressions are summarized afterwards. Perl-compatible regular expressions have different functionality, and are documented in the pcre2syntax(3) and pcre2pattern(3) manual pages, but work only if PCRE is available in the system.

3.1 Fundamental Structure

In regular expressions, the characters ‘.?*+{|()[\^$’ are special characters and have uses described below. All other characters are ordinary characters, and each ordinary character is a regular expression that matches itself.

The period ‘.’ matches any single character. It is unspecified whether ‘.’ matches an encoding error.

A regular expression may be followed by one of several repetition operators; the operators beginning with ‘{’ are called interval expressions.

‘?’ The preceding item is optional and is matched at most once.

‘*’ The preceding item is matched zero or more times.

‘+’ The preceding item is matched one or more times.

‘{n}’ The preceding item is matched exactly n times.

‘{n,}’ The preceding item is matched n or more times.

‘{,m}’ The preceding item is matched at most m times. This is a GNU extension.

‘{n,m}’ The preceding item is matched at least n times, but not more than m times.

The empty regular expression matches the empty string. Two regular expressions may be concatenated; the resulting regular expression matches any string formed by concatenating two substrings that respectively match the concatenated expressions.

Two regular expressions may be joined by the infix operator ‘|’. The resulting regular expression matches any string matching either of the two expressions, which are called alternatives.

Repetition takes precedence over concatenation, which in turn takes precedence over alternation. A whole expression may be enclosed in parentheses to override these precedence rules and form a subexpression. An unmatched ‘)’ matches just itself.

Not every character string is a valid regular expression. See Section 3.7 [Problematic Expressions], page 18.

3.2 Character Classes and Bracket Expressions

A bracket expression is a list of characters enclosed by ‘[’ and ‘]’. It matches any single character in that list. If the first character of the list is the caret ‘^’, then it matches any character not in the list, and it is unspecified whether it matches an encoding error. For example, the regular expression ‘[0123456789]’ matches any single digit, whereas ‘[^()]’ matches any single character that is not an opening or closing parenthesis, and might or might not match an encoding error.

Within a bracket expression, a range expression consists of two characters separated by a hyphen. It matches any single character that sorts between the two characters, inclusive. In the default C locale, the sorting sequence is the native character order; for example, ‘[a-d]’ is equivalent to ‘[abcd]’. In other locales, the sorting sequence is not specified, and ‘[a-d]’ might be equivalent to ‘[abcd]’ or to ‘[aBbCcDd]’, or it might fail to match any character, or the set of characters that it matches might be erratic, or it might be invalid. To obtain the traditional interpretation of bracket expressions, you can use the ‘C’ locale by setting the LC_ALL environment variable to the value ‘C’.

Finally, certain named classes of characters are predefined within bracket expressions, as follows. Their interpretation depends on the LC_CTYPE locale; for example, ‘[[:alnum:]]’ means the character class of numbers and letters in the current locale.

‘[:alnum:]’

Alphanumeric characters: ‘[:alpha:]’ and ‘[:digit:]’; in the ‘C’ locale and ASCII character encoding, this is the same as ‘[0-9A-Za-z]’.

‘[:alpha:]’

Alphabetic characters: ‘[:lower:]’ and ‘[:upper:]’; in the ‘C’ locale and ASCII character encoding, this is the same as ‘[A-Za-z]’.

‘[:blank:]’

Blank characters: space and tab.

‘[:cntrl:]’

Control characters. In ASCII, these characters have octal codes 000 through 037, and 177 (DEL). In other character sets, these are the equivalent characters, if any.

‘[:digit:]’
          Digits: 0 1 2 3 4 5 6 7 8 9.
‘[:graph:]’

Graphical characters: ‘[:alnum:]’ and ‘[:punct:]’.

‘[:lower:]’

Lower-case letters; in the ‘C’ locale and ASCII character encoding, this is a b c

          d e f g h i j k l m n o p q r s t u v w x y z.
‘[:print:]’

Printable characters: ‘[:alnum:]’, ‘[:punct:]’, and space.

‘[:punct:]’

Punctuation characters; in the ‘C’ locale and ASCII character encoding, this is

          ! " # $ % & ' ( ) * + , - . / : ; < = > ? @ [ \ ] ^_` { | } ~.

‘[:space:]’

Space characters: in the ‘C’ locale, this is tab, newline, vertical tab, form feed, carriage return, and space. See Chapter 4 [Usage], page 21, for more discussion of matching newlines.

‘[:upper:]’

Upper-case letters: in the ‘C’ locale and ASCII character encoding, this is A B

          C D E F G H I J K L M N O P Q R S T U V W X Y Z.
‘[:xdigit:]’
          Hexadecimal digits: 0 1 2 3 4 5 6 7 8 9 A B C D E F a b c d e f.

Note that the brackets in these class names are part of the symbolic names, and must be included in addition to the brackets delimiting the bracket expression.

If you mistakenly omit the outer brackets, and search for say, ‘[:upper:]’, GNU grep prints a diagnostic and exits with status 2, on the assumption that you did not intend to search for the regular expression ‘[:epru]’.

Special characters lose their special meaning inside bracket expressions. ‘]’ ends the bracket expression if it’s not the first list item. So, if you want to make

the ‘]’ character a list item, you must put it first.

‘[.’ represents the open collating symbol. ‘.]’ represents the close collating symbol. ‘[=’ represents the open equivalence class. ‘=]’ represents the close equivalence class. ‘[:’ represents the open character class symbol, and should be followed by a valid

character class name.

‘:]’ represents the close character class symbol. ‘-’ represents the range if it’s not first or last in a list or the ending point of a

range. To make the ‘-’ a list item, it is best to put it last.

‘^’ represents the characters not in the list. If you want to make the ‘^’ character a list item, place it anywhere but first.

3.3 Special Backslash Expressions

The ‘\’ character followed by a special character is a regular expression that matches the special character. The ‘\’ character, when followed by certain ordinary characters, takes a special meaning: ‘\b’ Match the empty string at the edge of a word. ‘\B’ Match the empty string provided it’s not at the edge of a word. ‘\<’ Match the empty string at the beginning of a word. ‘\>’ Match the empty string at the end of a word. ‘\w’ Match word constituent, it is a synonym for ‘[_[:alnum:]]’.

‘\W’ Match non-word constituent, it is a synonym for ‘[^_[:alnum:]]’.

‘\s’ Match whitespace, it is a synonym for ‘[[:space:]]’.

‘\S’ Match non-whitespace, it is a synonym for ‘[^[:space:]]’.

‘\]’ Match ‘]’.

‘\}’ Match ‘}’.

For example, ‘\brat\b’ matches the separate word ‘rat’, ‘\Brat\B’ matches ‘crate’ but

not ‘furry rat’.

The behavior of grep is unspecified if a unescaped backslash is not followed by a special character, a nonzero digit, or a character in the above list. Although grep might issue a diagnostic and/or give the backslash an interpretation now, its behavior may change if the syntax of regular expressions is extended in future versions.

3.4 Anchoring

The caret ‘^’ and the dollar sign ‘$’ are special characters that respectively match the empty string at the beginning and end of a line. They are termed anchors, since they force the match to be “anchored” to beginning or end of a line, respectively.

3.5 Back-references and Subexpressions

The back-reference ‘\n’, where n is a single nonzero digit, matches the substring previously matched by the nth parenthesized subexpression of the regular expression. For example, ‘(a)\1’ matches ‘aa’. If the parenthesized subexpression does not participate in the match, the back-reference makes the whole match fail; for example, ‘(a)*\1’ fails to match ‘a’. If the parenthesized subexpression matches more than one substring, the back-reference refers to the last matched substring; for example, ‘^(ab*)*\1$’ matches ‘ababbabb’ but not ‘ababbab’. When multiple regular expressions are given with -e or from a file (‘-f file’), back-references are local to each expression.

See Section 6.1 [Known Bugs], page 27, for some known problems with back-references.

3.6 Basic vs Extended Regular Expressions

Basic regular expressions differ from extended regular expressions in the following ways:

• The characters ‘?’, ‘+’, ‘{’, ‘|’, ‘(’, and ‘)’ lose their special meaning; instead use the backslashed versions ‘\?’, ‘\+’, ‘\{’, ‘\|’, ‘\(’, and ‘\)’. Also, a backslash is needed before an interval expression’s closing ‘}’.

• An unmatched ‘\)’ is invalid.

• If an unescaped ‘^’ appears neither first, nor directly after ‘\(’ or ‘\|’, it is treated like an ordinary character and is not an anchor.

• If an unescaped ‘$’ appears neither last, nor directly before ‘\|’ or ‘\)’, it is treated like an ordinary character and is not an anchor.

• If an unescaped ‘*’ appears first, or appears directly after ‘\(’ or ‘\|’ or anchoring ‘^’, it is treated like an ordinary character and is not a repetition operator.

3.7 Problematic Regular Expressions

Some strings are invalid regular expressions and cause grep to issue a diagnostic and fail. For example, ‘xy\1’ is invalid because there is no parenthesized subexpression for the backreference ‘\1’ to refer to.

Also, some regular expressions have unspecified behavior and should be avoided even if grep does not currently diagnose them. For example, ‘xy\0’ has unspecified behavior because ‘0’ is not a special character and ‘\0’ is not a special backslash expression (see Section 3.3 [Special Backslash Expressions], page 16). Unspecified behavior can be particularly problematic because the set of matched strings might be only partially specified, or not be specified at all, or the expression might even be invalid.

The following regular expression constructs are invalid on all platforms conforming to POSIX, so portable scripts can assume that grep rejects these constructs:

• A basic regular expression containing a back-reference ‘\n’ preceded by fewer than n closing parentheses. For example, ‘\(a\)\2’ is invalid.

• A bracket expression containing ‘[:’ that does not start a character class; and similarly for ‘[=’ and ‘[.’. For example, ‘[a[:b]’ and ‘[a[:ouch:]b]’ are invalid.

GNU grep treats the following constructs as invalid. However, other grep implementations might allow them, so portable scripts should not rely on their being invalid:

• Unescaped ‘\’ at the end of a regular expression.

• Unescaped ‘[’ that does not start a bracket expression.

• A ‘\{’ in a basic regular expression that does not start an interval expression.

• A basic regular expression with unbalanced ‘\(’ or ‘\)’, or an extended regular expression with unbalanced ‘(’.

• In the POSIX locale, a range expression like ‘z-a’ that represents zero elements. A non-GNU grep might treat it as a valid range that never matches.

• An interval expression with a repetition count greater than 32767. (The portable POSIX limit is 255, and even interval expressions with smaller counts can be impractically slow on all known implementations.)

• A bracket expression that contains at least three elements, the first and last of which are both ‘:’, or both ‘.’, or both ‘=’. For example, a non-GNU grep might treat

  ‘[:alpha:]’ like ‘[[:alpha:]]’, or like ‘[:ahlp]’.

The following constructs have well-defined behavior in GNU grep. However, they have unspecified behavior elsewhere, so portable scripts should avoid them:

• Special backslash expressions like ‘\b’, ‘\<’, and ‘\]’. See Section 3.3 [Special Backslash Expressions], page 16.

• A basic regular expression that uses ‘\?’, ‘\+’, or ‘\|’.

• An extended regular expression that uses back-references.

• An empty regular expression, subexpression, or alternative. For example, ‘(a|bc|)’ is not portable; a portable equivalent is ‘(a|bc)?’.

• In a basic regular expression, an anchoring ‘^’ that appears directly after ‘\(’, or an anchoring ‘$’ that appears directly before ‘\)’.

• In a basic regular expression, a repetition operator that directly follows another repetition operator.

• In an extended regular expression, unescaped ‘{’ that does not begin a valid interval expression. GNU grep treats the ‘{’ as an ordinary character.

• A null character or an encoding error in either pattern or input data. See Section 3.8 [Character Encoding], page 19.

• An input file that ends in a non-newline character, where GNU grep silently supplies a newline.

The following constructs have unspecified behavior, in both GNU and other grep implementations. Scripts should avoid them whenever possible.

• A backslash escaping an ordinary character, unless it is a back-reference like ‘\1’ or a special backslash expression like ‘\<’ or ‘\b’. See Section 3.3 [Special Backslash Expressions], page 16. For example, ‘\x’ has unspecified behavior now, and a future version of grep might specify ‘\x’ to have a new behavior.

• A repetition operator that appears directly after an anchor, or at the start of a complete regular expression, parenthesized subexpression, or alternative. For example, ‘+|^*(+a|?-b)’ has unspecified behavior, whereas ‘\+|^\*(\+a|\?-b)’ is portable.

• A range expression outside the POSIX locale. For example, in some locales ‘[a-z]’ might match some characters that are not lowercase letters, or might not match some lowercase letters, or might be invalid. With GNU grep it is not documented whether these range expressions use native code points, or use the collating sequence specified by the LC_COLLATE category, or have some other interpretation. Outside the POSIX locale, it is portable to use ‘[[:lower:]]’ to match a lower-case letter, or ‘[abcdefghijklmnopqrstuvwxyz]’ to match an ASCII lower-case letter.

3.8 Character Encoding

The LC_CTYPE locale specifies the encoding of characters in patterns and data, that is, whether text is encoded in UTF-8, ASCII, or some other encoding. See Section 2.2 [Environment Variables], page 9.

In the ‘C’ or ‘POSIX’ locale, every character is encoded as a single byte and every byte is a valid character. In more-complex encodings such as UTF-8, a sequence of multiple bytes may be needed to represent a character, and some bytes may be encoding errors that do not contribute to the representation of any character. POSIX does not specify the behavior of grep when patterns or input data contain encoding errors or null characters, so portable scripts should avoid such usage. As an extension to POSIX, GNU grep treats null characters like any other character. However, unless the -a (--binary-files=text) option is used, the presence of null characters in input or of encoding errors in output causes GNU grep to treat the file as binary and suppress details about matches. See Section 2.1.6 [File and Directory Selection], page 7.

Regardless of locale, the 103 characters in the POSIX Portable Character Set (a subset of ASCII) are always encoded as a single byte, and the 128 ASCII characters have their usual single-byte encodings on all but oddball platforms.

3.9 Matching Non-ASCII and Non-printable Characters

In a regular expression, non-ASCII and non-printable characters other than newline are not special, and represent themselves. For example, in a locale using UTF-8 the command ‘grep 'Λ ω'’ (where the white space between ‘Λ’ and the ‘ω’ is a tab character) searches for ‘Λ’ (Unicode character U+039B GREEK CAPITAL LETTER LAMBDA), followed by a tab (U+0009 TAB), followed by ‘ω’ (U+03C9 GREEK SMALL LETTER OMEGA).

Suppose you want to limit your pattern to only printable characters (or even only printable ASCII characters) to keep your script readable or portable, but you also want to match specific non-ASCII or non-null non-printable characters. If you are using the -P (--perl-regexp) option, PCREs give you several ways to do this. Otherwise, if you are using Bash, the GNU project’s shell, you can represent these characters via ANSI-C quoting. For example, the Bash commands ‘grep $'Λ\tω'’ and ‘grep $'\u039B\t\u03C9'’ both search for the same three-character string ‘Λ ω’ mentioned earlier. However, because Bash translates ANSI-C quoting before grep sees the pattern, this technique should not be used to match printable ASCII characters; for example, ‘grep $'\u005E'’ is equivalent to ‘grep '^'’ and matches any line, not just lines containing the character ‘^’ (U+005E CIRCUMFLEX AC-CENT).

Since PCREs and ANSI-C quoting are GNU extensions to POSIX, portable shell scripts written in ASCII should use other methods to match specific non-ASCII characters. For

example, in a UTF-8 locale the command ‘grep "$(printf '\316\233\t\317\211\n')"’

is a portable albeit hard-to-read alternative to Bash’s ‘grep $'Λ\tω'’. However, none of these techniques will let you put a null character directly into a command-line pattern; null characters can appear only in a pattern specified via the -f (--file) option.

4 Usage

Here is an example command that invokes GNU grep:

     grep -i 'hello.*world' menu.h main.c

This lists all lines in the files menu.h and main.c that contain the string ‘hello’ followed by the string ‘world’; this is because ‘.*’ matches zero or more characters within a line. See Chapter 3 [Regular Expressions], page 14. The -i option causes grep to ignore case, causing it to match the line ‘Hello, world!’, which it would not otherwise match.

Here is a more complex example, showing the location and contents of any line containing ‘f’ and ending in ‘.c’, within all files in the current directory whose names start with non- ‘.’, contain ‘g’, and end in ‘.h’. The -n option outputs line numbers, the -- argument treats any later arguments as file names not options even if *g*.h expands to a file name that starts with ‘-’, and the empty file /dev/null causes file names to be output even if only one file name happens to be of the form ‘*g*.h’.

     grep -n -- 'f.*\.c$' *g*.h /dev/null

Note that the regular expression syntax used in the pattern differs from the globbing syntax that the shell uses to match file names.

See Chapter 2 [Invoking], page 2, for more details about how to invoke grep.

Here are some common questions and answers about grep usage.

1. How can I list just the names of matching files?

         grep -l 'main' test-*.c

lists names of ‘test-*.c’ files in the current directory whose contents mention ‘main’.

2. How do I search directories recursively?

         grep -r 'hello' /home/gigi

searches for ‘hello’ in all files under the /home/gigi directory. For more control over which files are searched, use find and grep. For example, the following command searches only C files:

         find /home/gigi -name '*.c' ! -type d \
           -exec grep -H 'hello' '{}' +

This differs from the command:

         grep -H 'hello' /home/gigi/*.c

which merely looks for ‘hello’ in non-hidden C files in /home/gigi whose names end in ‘.c’. The find command line above is more similar to the command:

         grep -r --include='*.c' 'hello' /home/gigi

3. What if a pattern or file has a leading ‘-’? For example:

         grep "$pattern" *

can behave unexpectedly if the value of ‘pattern’ begins with ‘-’, or if the ‘*’ expands to a file name with leading ‘-’. To avoid the problem, you can use -e for patterns and leading ‘./’ for files:

         grep -e "$pattern" ./*

searches for all lines matching the pattern in all the working directory’s files whose names do not begin with ‘.’. Without the -e, grep might treat the pattern as an option if it begins with ‘-’. Without the ‘./’, there might be similar problems with file names beginning with ‘-’. Alternatively, you can use ‘--’ before the pattern and file names:

     grep -- "$pattern" *

This also fixes the problem, except that if there is a file named ‘-’, grep misinterprets the ‘-’ as standard input.

4. Suppose I want to search for a whole word, not a part of a word?

     grep -w 'hello' test*.log

searches only for instances of ‘hello’ that are entire words; it does not match ‘Othello’. For more control, use ‘\<’ and ‘\>’ to match the start and end of words. For example:

     grep 'hello\>' test*.log

searches only for words ending in ‘hello’, so it matches the word ‘Othello’.

5. How do I output context around the matching lines?

     grep -C 2 'hello' test*.log

prints two lines of context around each matching line.

6. How do I force grep to print the name of the file?

Append /dev/null:
     grep 'eli' /etc/passwd /dev/null

gets you:

     /etc/passwd:eli:x:2098:1000:Eli Smith:/home/eli:/bin/bash

Alternatively, use -H, which is a GNU extension:

     grep -H 'eli' /etc/passwd

7. Why do people use strange regular expressions on ps output?

     ps -ef | grep '[c]ron'

If the pattern had been written without the square brackets, it would have matched not only the ps output line for cron, but also the ps output line for grep. Note that on some platforms, ps limits the output to the width of the screen; grep does not have any limit on the length of a line except the available memory.

8. Why does grep report “Binary file matches”?

If grep listed all matching “lines” from a binary file, it would probably generate output that is not useful, and it might even muck up your display. So GNU grep suppresses output from files that appear to be binary files. To force GNU grep to output lines even from files that appear to be binary, use the -a or ‘--binary-files=text’ option. To eliminate the “Binary file matches” messages, use

the -I or ‘--binary-files=without-match’ option.

9. Why doesn’t ‘grep -lv’ print non-matching file names? ‘grep -lv’ lists the names of all files containing one or more lines that do not match. To list the names of all files that contain no matching lines, use the -L or --files-

without-match option.

10. I can do “OR” with ‘|’, but what about “AND”?

     grep 'paul' /etc/motd | grep 'franc,ois'

finds all lines that contain both ‘paul’ and ‘franc,ois’.

11. Why does the empty pattern match every input line? The grep command searches for lines that contain strings that match a pattern. Every line contains the empty string, so an empty pattern causes grep to find a match on each line. It is not the only such pattern: ‘^’, ‘$’, and many other patterns cause grep to match every line. To match empty lines, use the pattern ‘^$’. To match blank lines, use the pattern ‘^[[:blank:]]*$’. To match no lines at all, use an extended regular expression like ‘a^’ or ‘$a’. To match every line, a portable script should use a pattern like ‘^’ instead of the empty pattern, as POSIX does not specify the behavior of the empty pattern.

12. How can I search in both standard input and in files?

Use the special file name ‘-’:

     cat /etc/passwd | grep 'alain' - /etc/motd

13. Why can’t I combine the shell’s ‘set -e’ with grep? The grep command follows the convention of programs like cmp and diff where an exit status of 1 is not an error. The shell command ‘set -e’ causes the shell to exit if any subcommand exits with nonzero status, and this will cause the shell to exit merely because grep selected no lines, which is ordinarily not what you want. There is a related problem with Bash’s set -e -o pipefail. Since grep does not always read all its input, a command outputting to a pipe read by grep can fail when grep exits before reading all its input, and the command’s failure can cause Bash to exit.

14. Why is this back-reference failing?

     echo 'ba' | grep -E '(a)\1|b\1'

This outputs an error message, because the second ‘\1’ has nothing to refer back to, meaning it will never match anything.

15. How can I match across lines?

Standard grep cannot do this, as it is fundamentally line-based. Therefore, merely using the [:space:] character class does not match newlines in the way you might expect. With the GNU grep option -z (--null-data), each input and output “line” is nullterminated; see Section 2.1.7 [Other Options], page 9. Thus, you can match newlines in the input, but typically if there is a match the entire input is output, so this usage is often combined with output-suppressing options like -q, e.g.:

     printf 'foo\nbar\n' | grep -z -q 'foo[[:space:]]\+bar'

If this does not suffice, you can transform the input before giving it to grep, or turn to awk, sed, perl, or many other utilities that are designed to operate across lines.

16. What do grep, -E, and -F stand for?

The name grep comes from the way line editing was done on Unix. For example, ed uses the following syntax to print a list of matching lines on the screen:

     global/regular expression/print
     g/re/p

The -E option stands for Extended grep. The -F option stands for Fixed grep; 17. What happened to egrep and fgrep?

7th Edition Unix had commands egrep and fgrep that were the counterparts of the modern ‘grep -E’ and ‘grep -F’. Although breaking up grep into three programs was perhaps useful on the small computers of the 1970s, egrep and fgrep were deemed obsolescent by POSIX in 1992, removed from POSIX in 2001, deprecated by GNU Grep 2.5.3 in 2007, and changed to issue obsolescence warnings by GNU Grep 3.8 in 2022; eventually, they are planned to be removed entirely. If you prefer the old names, you can use your own substitutes, such as a shell script named egrep with the following contents:

     #!/bin/sh
     exec grep -E "$@"

5 Performance

Typically grep is an efficient way to search text. However, it can be quite slow in some cases, and it can search large files where even minor performance tweaking can help significantly. Although the algorithm used by grep is an implementation detail that can change from release to release, understanding its basic strengths and weaknesses can help you improve its performance.

The grep command operates partly via a set of automata that are designed for efficiency, and partly via a slower matcher that takes over when the fast matchers run into unusual features like back-references. When feasible, the Boyer–Moore fast string searching algorithm is used to match a single fixed pattern, and the Aho–Corasick algorithm is used to match multiple fixed patterns.

Generally speaking grep operates more efficiently in single-byte locales, since it can avoid the special processing needed for multi-byte characters. If your patterns will work just as well that way, setting LC_ALL to a single-byte locale can help performance considerably. Setting ‘LC_ALL='C'’ can be particularly efficient, as grep is tuned for that locale.

Outside the ‘C’ locale, case-insensitive search, and search for bracket expressions like ‘[a-z]’ and ‘[[=a=]b]’, can be surprisingly inefficient due to difficulties in fast portable access to concepts like multi-character collating elements.

Interval expressions may be implemented internally via repetition. For example,

‘^(a|bc){2,4}$’ might be implemented as ‘^(a|bc)(a|bc)((a|bc)(a|bc)?)?$’. A large

repetition count may exhaust memory or greatly slow matching. Even small counts can cause problems if cascaded; for example, ‘grep -E ".*{10,}{10,}{10,}{10,}{10,}"’ is likely to overflow a stack. Fortunately, regular expressions like these are typically artificial, and cascaded repetitions do not conform to POSIX so cannot be used in portable programs anyway.

A back-reference such as ‘\1’ can hurt performance significantly in some cases, since back-references cannot in general be implemented via a finite state automaton, and instead trigger a backtracking algorithm that can be quite inefficient. For example, although the pattern ‘^(.*)\1{14}(.*)\2{13}$’ matches only lines whose lengths can be written as a sum 15x + 14y for nonnegative integers x and y, the pattern matcher does not perform linear Diophantine analysis and instead backtracks through all possible matching strings, using an algorithm that is exponential in the worst case.

On some operating systems that support files with holes—large regions of zeros that are not physically present on secondary storage—grep can skip over the holes efficiently without needing to read the zeros. This optimization is not available if the -a (--binary-files=text) option is used (see Section 2.1.6 [File and Directory Selection], page 7), unless the -z (--null-data) option is also used (see Section 2.1.7 [Other Options], page 9).

For efficiency grep does not always read all its input. For example, the shell command ‘sed '/^...$/d' | grep -q X’ can cause grep to exit immediately after reading a line containing ‘X’, without bothering to read the rest of its input data. This in turn can cause sed to exit with a nonzero status because sed cannot write to its output pipe after grep exits.

For more about the algorithms used by grep and about related string matching algorithms, see:

• Aho AV. Algorithms for finding patterns in strings. In: van Leeuwen J. Handbook of Theoretical Computer Science, vol. A. New York: Elsevier; 1990. p. 255–300. This surveys classic string matching algorithms, some of which are used by grep.

• Aho AV, Corasick MJ. Efficient string matching: an aid to bibliographic search. CACM.

1975;18(6):333–40. https://doi.org/10.1145/360825.360855. This introduces the

Aho–Corasick algorithm.

• Boyer RS, Moore JS. A fast string searching algorithm. CACM. 1977;20(10):762–72.

https://doi.org/10.1145/359842.359859. This introduces the Boyer–Moore algo-

rithm.

• Faro S, Lecroq T. The exact online string matching problem: a review of the most recent results. ACM Comput Surv. 2013;45(2):13. https://doi.org/10.1145/2431211. 2431212. This surveys string matching algorithms that might help improve the performance of grep in the future.

• Hakak SI, Kamsin A, Shivakumara P, Gilkar GA, Khan WZ, Imran M. Exact string matching algorithms: survey issues, and future research directions. IEEE Access.

2019;7:69614–37. https://doi.org/10.1109/ACCESS.2019.2914071. This survey is

more recent than Faro & Lecroq, and focuses on taxonomy instead of performance.

• Hume A, Sunday D. Fast string search. Software Pract Exper. 1991;21(11):1221–48.

https://doi.org/10.1002/spe.4380211105. This excellent albeit now-dated survey

aided the initial development of grep.

6 Reporting bugs

Bug reports can be found at the GNU bug report logs for grep (https://debbugs.gnu. org/cgi/pkgreport.cgi?package=grep). If you find a bug not listed there, please email it to bug-grep@gnu.org to create a new bug report.

6.1 Known Bugs

Large repetition counts in the ‘{n,m}’ construct may cause grep to use lots of memory. In addition, certain other obscure regular expressions require exponential time and space, and may cause grep to run out of memory.

Back-references can greatly slow down matching, as they can generate exponentially many matching possibilities that can consume both time and memory to explore. Also, the POSIX specification for back-references is at times unclear. Furthermore, many regular expression implementations have back-reference bugs that can cause programs to return incorrect answers or even crash, and fixing these bugs has often been low-priority: for example, as of 2021 the GNU C library bug database (https://sourceware.org/bugzilla/ ) contained back-reference bugs 52, 10844, 11053, 24269 and 25322, with little sign of forthcoming fixes. Luckily, back-references are rarely useful and it should be little trouble to avoid them in practical applications.

7 Copying

GNU grep is licensed under the GNU GPL, which makes it free software.

The “free” in “free software” refers to liberty, not price. As some GNU project advocates like to point out, think of “free speech” rather than “free beer”. In short, you have the right (freedom) to run and change grep and distribute it to other people, and—if you want— charge money for doing either. The important restriction is that you have to grant your recipients the same rights and impose the same restrictions.

This general method of licensing software is sometimes called open source. The GNU project prefers the term “free software” for reasons outlined at https://www.gnu.org/

philosophy/open-source-misses-the-point.html.

This manual is free documentation in the same sense. The documentation license is included below. The license for the program is available with the source code, or at https://

www.gnu.org/licenses/gpl.html.

7.1 GNU Free Documentation License

Version 1.3, 3 November 2008

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2000–2002, 2007–2008, 2023 Free Software Foundation, Inc.

 https://fsf.org/

Everyone is permitted to copy and distribute verbatim copies of this license document, but changing it is not allowed.

0. PREAMBLE The purpose of this License is to make a manual, textbook, or other functional and useful document free in the sense of freedom: to assure everyone the effective freedom to copy and redistribute it, with or without modifying it, either commercially or noncommercially. Secondarily, this License preserves for the author and publisher a way to get credit for their work, while not being considered responsible for modifications made by others. This License is a kind of “copyleft”, which means that derivative works of the document must themselves be free in the same sense. It complements the GNU General Public License, which is a copyleft license designed for free software. We have designed this License in order to use it for manuals for free software, because free software needs free documentation: a free program should come with manuals providing the same freedoms that the software does. But this License is not limited to software manuals; it can be used for any textual work, regardless of subject matter or whether it is published as a printed book. We recommend this License principally for works whose purpose is instruction or reference.

1. APPLICABILITY AND DEFINITIONS This License applies to any manual or other work, in any medium, that contains a notice placed by the copyright holder saying it can be distributed under the terms of this License. Such a notice grants a world-wide, royalty-free license, unlimited in duration, to use that work under the conditions stated herein. The “Document”, below, refers to any such manual or work. Any member of the public is a licensee, and is addressed as “you”. You accept the license if you copy, modify or distribute the work in a way requiring permission under copyright law.

A “Modified Version” of the Document means any work containing the Document or a portion of it, either copied verbatim, or with modifications and/or translated into another language.

A “Secondary Section” is a named appendix or a front-matter section of the Document that deals exclusively with the relationship of the publishers or authors of the Document to the Document’s overall subject (or to related matters) and contains nothing that could fall directly within that overall subject. (Thus, if the Document is in part a textbook of mathematics, a Secondary Section may not explain any mathematics.) The relationship could be a matter of historical connection with the subject or with related matters, or of legal, commercial, philosophical, ethical or political position regarding them.

The “Invariant Sections” are certain Secondary Sections whose titles are designated, as being those of Invariant Sections, in the notice that says that the Document is released under this License. If a section does not fit the above definition of Secondary then it is not allowed to be designated as Invariant. The Document may contain zero Invariant Sections. If the Document does not identify any Invariant Sections then there are none.

The “Cover Texts” are certain short passages of text that are listed, as Front-Cover Texts or Back-Cover Texts, in the notice that says that the Document is released under this License. A Front-Cover Text may be at most 5 words, and a Back-Cover Text may be at most 25 words.

A “Transparent” copy of the Document means a machine-readable copy, represented in a format whose specification is available to the general public, that is suitable for revising the document straightforwardly with generic text editors or (for images composed of pixels) generic paint programs or (for drawings) some widely available drawing editor, and that is suitable for input to text formatters or for automatic translation to a variety of formats suitable for input to text formatters. A copy made in an otherwise Transparent file format whose markup, or absence of markup, has been arranged to thwart or discourage subsequent modification by readers is not Transparent. An image format is not Transparent if used for any substantial amount of text. A copy that is not “Transparent” is called “Opaque”.

Examples of suitable formats for Transparent copies include plain ASCII without markup, Texinfo input format, LaTEX input format, SGML or XML using a publicly available DTD, and standard-conforming simple HTML, PostScript or PDF designed for human modification. Examples of transparent image formats include PNG, XCF and JPG. Opaque formats include proprietary formats that can be read and edited only by proprietary word processors, SGML or XML for which the DTD and/or processing tools are not generally available, and the machine-generated HTML, PostScript or PDF produced by some word processors for output purposes only.

The “Title Page” means, for a printed book, the title page itself, plus such following pages as are needed to hold, legibly, the material this License requires to appear in the title page. For works in formats which do not have any title page as such, “Title Page” means the text near the most prominent appearance of the work’s title, preceding the beginning of the body of the text. The “publisher” means any person or entity that distributes copies of the Document to the public. A section “Entitled XYZ” means a named subunit of the Document whose title either is precisely XYZ or contains XYZ in parentheses following text that translates XYZ in another language. (Here XYZ stands for a specific section name mentioned below, such as “Acknowledgements”, “Dedications”, “Endorsements”, or “History”.) To “Preserve the Title” of such a section when you modify the Document means that it remains a section “Entitled XYZ” according to this definition. The Document may include Warranty Disclaimers next to the notice which states that this License applies to the Document. These Warranty Disclaimers are considered to be included by reference in this License, but only as regards disclaiming warranties: any other implication that these Warranty Disclaimers may have is void and has no effect on the meaning of this License.

2. VERBATIM COPYING You may copy and distribute the Document in any medium, either commercially or noncommercially, provided that this License, the copyright notices, and the license notice saying this License applies to the Document are reproduced in all copies, and that you add no other conditions whatsoever to those of this License. You may not use technical measures to obstruct or control the reading or further copying of the copies you make or distribute. However, you may accept compensation in exchange for copies. If you distribute a large enough number of copies you must also follow the conditions in section 3. You may also lend copies, under the same conditions stated above, and you may publicly display copies.

3. COPYING IN QUANTITY If you publish printed copies (or copies in media that commonly have printed covers) of the Document, numbering more than 100, and the Document’s license notice requires Cover Texts, you must enclose the copies in covers that carry, clearly and legibly, all these Cover Texts: Front-Cover Texts on the front cover, and Back-Cover Texts on the back cover. Both covers must also clearly and legibly identify you as the publisher of these copies. The front cover must present the full title with all words of the title equally prominent and visible. You may add other material on the covers in addition. Copying with changes limited to the covers, as long as they preserve the title of the Document and satisfy these conditions, can be treated as verbatim copying in other respects. If the required texts for either cover are too voluminous to fit legibly, you should put the first ones listed (as many as fit reasonably) on the actual cover, and continue the rest onto adjacent pages. If you publish or distribute Opaque copies of the Document numbering more than 100, you must either include a machine-readable Transparent copy along with each Opaque copy, or state in or with each Opaque copy a computer-network location from which the general network-using public has access to download using public-standard network protocols a complete Transparent copy of the Document, free of added material. If you use the latter option, you must take reasonably prudent steps, when you begin distribution of Opaque copies in quantity, to ensure that this Transparent copy will remain thus accessible at the stated location until at least one year after the last time you distribute an Opaque copy (directly or through your agents or retailers) of that edition to the public. It is requested, but not required, that you contact the authors of the Document well before redistributing any large number of copies, to give them a chance to provide you with an updated version of the Document.

4. MODIFICATIONS You may copy and distribute a Modified Version of the Document under the conditions of sections 2 and 3 above, provided that you release the Modified Version under precisely this License, with the Modified Version filling the role of the Document, thus licensing distribution and modification of the Modified Version to whoever possesses a copy of it. In addition, you must do these things in the Modified Version: A. Use in the Title Page (and on the covers, if any) a title distinct from that of the

Document, and from those of previous versions (which should, if there were any, be listed in the History section of the Document). You may use the same title as a previous version if the original publisher of that version gives permission.

B. List on the Title Page, as authors, one or more persons or entities responsible for authorship of the modifications in the Modified Version, together with at least five of the principal authors of the Document (all of its principal authors, if it has fewer than five), unless they release you from this requirement.

C. State on the Title page the name of the publisher of the Modified Version, as the

publisher. D. Preserve all the copyright notices of the Document. E. Add an appropriate copyright notice for your modifications adjacent to the other

copyright notices. F. Include, immediately after the copyright notices, a license notice giving the public

permission to use the Modified Version under the terms of this License, in the form shown in the Addendum below.

G. Preserve in that license notice the full lists of Invariant Sections and required Cover

Texts given in the Document’s license notice. H. Include an unaltered copy of this License. I. Preserve the section Entitled “History”, Preserve its Title, and add to it an item

stating at least the title, year, new authors, and publisher of the Modified Version as given on the Title Page. If there is no section Entitled “History” in the Document, create one stating the title, year, authors, and publisher of the Document as given on its Title Page, then add an item describing the Modified Version as stated in the previous sentence.

J. Preserve the network location, if any, given in the Document for public access to a Transparent copy of the Document, and likewise the network locations given in the Document for previous versions it was based on. These may be placed in the “History” section. You may omit a network location for a work that was published at least four years before the Document itself, or if the original publisher of the version it refers to gives permission.

K. For any section Entitled “Acknowledgements” or “Dedications”, Preserve the Title of the section, and preserve in the section all the substance and tone of each of the contributor acknowledgements and/or dedications given therein.

L. Preserve all the Invariant Sections of the Document, unaltered in their text and in their titles. Section numbers or the equivalent are not considered part of the section titles.

M. Delete any section Entitled “Endorsements”. Such a section may not be included

in the Modified Version. N. Do not retitle any existing section to be Entitled “Endorsements” or to conflict in

title with any Invariant Section. O. Preserve any Warranty Disclaimers.

If the Modified Version includes new front-matter sections or appendices that qualify as Secondary Sections and contain no material copied from the Document, you may at your option designate some or all of these sections as invariant. To do this, add their titles to the list of Invariant Sections in the Modified Version’s license notice. These titles must be distinct from any other section titles. You may add a section Entitled “Endorsements”, provided it contains nothing but endorsements of your Modified Version by various parties—for example, statements of peer review or that the text has been approved by an organization as the authoritative definition of a standard. You may add a passage of up to five words as a Front-Cover Text, and a passage of up to 25 words as a Back-Cover Text, to the end of the list of Cover Texts in the Modified Version. Only one passage of Front-Cover Text and one of Back-Cover Text may be added by (or through arrangements made by) any one entity. If the Document already includes a cover text for the same cover, previously added by you or by arrangement made by the same entity you are acting on behalf of, you may not add another; but you may replace the old one, on explicit permission from the previous publisher that added the old one. The author(s) and publisher(s) of the Document do not by this License give permission to use their names for publicity for or to assert or imply endorsement of any Modified Version.

5. COMBINING DOCUMENTS You may combine the Document with other documents released under this License, under the terms defined in section 4 above for modified versions, provided that you include in the combination all of the Invariant Sections of all of the original documents, unmodified, and list them all as Invariant Sections of your combined work in its license notice, and that you preserve all their Warranty Disclaimers. The combined work need only contain one copy of this License, and multiple identical Invariant Sections may be replaced with a single copy. If there are multiple Invariant Sections with the same name but different contents, make the title of each such section unique by adding at the end of it, in parentheses, the name of the original author or publisher of that section if known, or else a unique number. Make the same adjustment to the section titles in the list of Invariant Sections in the license notice of the combined work.

In the combination, you must combine any sections Entitled “History” in the various original documents, forming one section Entitled “History”; likewise combine any sections Entitled “Acknowledgements”, and any sections Entitled “Dedications”. You must delete all sections Entitled “Endorsements.”

6. COLLECTIONS OF DOCUMENTS You may make a collection consisting of the Document and other documents released under this License, and replace the individual copies of this License in the various documents with a single copy that is included in the collection, provided that you follow the rules of this License for verbatim copying of each of the documents in all other respects. You may extract a single document from such a collection, and distribute it individually under this License, provided you insert a copy of this License into the extracted document, and follow this License in all other respects regarding verbatim copying of that document.

7. AGGREGATION WITH INDEPENDENT WORKS A compilation of the Document or its derivatives with other separate and independent documents or works, in or on a volume of a storage or distribution medium, is called an “aggregate” if the copyright resulting from the compilation is not used to limit the legal rights of the compilation’s users beyond what the individual works permit. When the Document is included in an aggregate, this License does not apply to the other works in the aggregate which are not themselves derivative works of the Document. If the Cover Text requirement of section 3 is applicable to these copies of the Document, then if the Document is less than one half of the entire aggregate, the Document’s Cover Texts may be placed on covers that bracket the Document within the aggregate, or the electronic equivalent of covers if the Document is in electronic form. Otherwise they must appear on printed covers that bracket the whole aggregate.

8. TRANSLATION Translation is considered a kind of modification, so you may distribute translations of the Document under the terms of section 4. Replacing Invariant Sections with translations requires special permission from their copyright holders, but you may include translations of some or all Invariant Sections in addition to the original versions of these Invariant Sections. You may include a translation of this License, and all the license notices in the Document, and any Warranty Disclaimers, provided that you also include the original English version of this License and the original versions of those notices and disclaimers. In case of a disagreement between the translation and the original version of this License or a notice or disclaimer, the original version will prevail. If a section in the Document is Entitled “Acknowledgements”, “Dedications”, or “History”, the requirement (section 4) to Preserve its Title (section 1) will typically require changing the actual title.

9. TERMINATION You may not copy, modify, sublicense, or distribute the Document except as expressly provided under this License. Any attempt otherwise to copy, modify, sublicense, or distribute it is void, and will automatically terminate your rights under this License.

However, if you cease all violation of this License, then your license from a particular copyright holder is reinstated (a) provisionally, unless and until the copyright holder explicitly and finally terminates your license, and (b) permanently, if the copyright holder fails to notify you of the violation by some reasonable means prior to 60 days after the cessation. Moreover, your license from a particular copyright holder is reinstated permanently if the copyright holder notifies you of the violation by some reasonable means, this is the first time you have received notice of violation of this License (for any work) from that copyright holder, and you cure the violation prior to 30 days after your receipt of the notice. Termination of your rights under this section does not terminate the licenses of parties who have received copies or rights from you under this License. If your rights have been terminated and not permanently reinstated, receipt of a copy of some or all of the same material does not give you any rights to use it.

10. FUTURE REVISIONS OF THIS LICENSE The Free Software Foundation may publish new, revised versions of the GNU Free Documentation License from time to time. Such new versions will be similar in spirit to the present version, but may differ in detail to address new problems or concerns.

See https://www.gnu.org/licenses/.

Each version of the License is given a distinguishing version number. If the Document specifies that a particular numbered version of this License “or any later version” applies to it, you have the option of following the terms and conditions either of that specified version or of any later version that has been published (not as a draft) by the Free Software Foundation. If the Document does not specify a version number of this License, you may choose any version ever published (not as a draft) by the Free Software Foundation. If the Document specifies that a proxy can decide which future versions of this License can be used, that proxy’s public statement of acceptance of a version permanently authorizes you to choose that version for the Document.

11. RELICENSING “Massive Multiauthor Collaboration Site” (or “MMC Site”) means any World Wide Web server that publishes copyrightable works and also provides prominent facilities for anybody to edit those works. A public wiki that anybody can edit is an example of such a server. A “Massive Multiauthor Collaboration” (or “MMC”) contained in the site means any set of copyrightable works thus published on the MMC site. “CC-BY-SA” means the Creative Commons Attribution-Share Alike 3.0 license published by Creative Commons Corporation, a not-for-profit corporation with a principal place of business in San Francisco, California, as well as future copyleft versions of that license published by that same organization. “Incorporate” means to publish or republish a Document, in whole or in part, as part of another Document. An MMC is “eligible for relicensing” if it is licensed under this License, and if all works that were first published under this License somewhere other than this MMC, and subsequently incorporated in whole or in part into the MMC, (1) had no cover texts or invariant sections, and (2) were thus incorporated prior to November 1, 2008.

The operator of an MMC Site may republish an MMC contained in the site under CC-BY-SA on the same site at any time before August 1, 2009, provided the MMC is eligible for relicensing.

ADDENDUM: How to use this License for your documents

To use this License in a document you have written, include a copy of the License in the document and put the following copyright and license notices just after the title page:

        Copyright (C)  year  your name.
        Permission is granted to copy, distribute and/or modify this document
        under the terms of the GNU Free Documentation License, Version 1.3
        or any later version published by the Free Software Foundation;
        with no Invariant Sections, no Front-Cover Texts, and no Back-Cover
        Texts.  A copy of the license is included in the section entitled `  ` GNU
        Free Documentation License''.

If you have Invariant Sections, Front-Cover Texts and Back-Cover Texts, replace the “with…Texts.” line with this:

          with the Invariant Sections being list their titles, with
          the Front-Cover Texts being list, and with the Back-Cover Texts
          being list.

If you have Invariant Sections without Cover Texts, or some other combination of the three, merge those two alternatives to suit the situation.

If your document contains nontrivial examples of program code, we recommend releasing these examples in parallel under your choice of free software license, such as the GNU General Public License, to permit their use in free software.

Index

*

* . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

+

+ . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

–

-- . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 --after-context . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 --basic-regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 --before-context . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 --binary . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 --binary-files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 --byte-offset. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 --color . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 --colour . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 --context . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 --count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

--dereference-recursive . . . . . . . . . . . . . . . . . . . . . . 9

--devices . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 --directories. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 --exclude . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 --exclude-dir. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 --exclude-from . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 --extended-regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 --file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 --files-with-matches. . . . . . . . . . . . . . . . . . . . . . . . . . 4 --files-without-match . . . . . . . . . . . . . . . . . . . . . . . . 4 --fixed-strings . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 --group-separator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 --help. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 --ignore-case. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 --include . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 --initial-tab. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 --invert-match . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 --label . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 --line-buffered . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 --line-number. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 --line-regexp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 --max-count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 --no-filename. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 --no-ignore-case . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 --no-messages. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 --null. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 --null-data . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 --only-matching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 --perl-regexp . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 --quiet . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 --recursive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 --regexp=patterns . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 --silent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 --text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 --version . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 --with-filename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 --word-regexp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 -a . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 -A . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 -b . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 -B . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 -c . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 -C . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 -d . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 -D . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 -e . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 -E . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 -f . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 -F . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 -G . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 -h . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 -H . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 -i . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 -l . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 -L . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 -m . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 -n . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 -num . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 -o . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 -P . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 -q . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 -r . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 -R . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 -s . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 -T . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 -U . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 -v . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 -V . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 -w . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 -x . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 -y . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 -z . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 -Z . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

.

. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

?

? . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

{

{,m} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 {n,} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 {n,m}. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 {n} . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

Index 38

A

after context . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 alnum character class. . . . . . . . . . . . . . . . . . . . . . . . . . . 15 alpha character class. . . . . . . . . . . . . . . . . . . . . . . . . . . 15 alphabetic characters . . . . . . . . . . . . . . . . . . . . . . . . . . 15 alphanumeric characters . . . . . . . . . . . . . . . . . . . . . . . 15 alternatives in regular expressions . . . . . . . . . . . . . . 14 anchoring . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 asterisk . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

B

back-reference . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 back-references. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 backslash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 basic regular expressions . . . . . . . . . . . . . . . . . . . . . . . 17 before context. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 binary files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 binary I/O . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 blank character class. . . . . . . . . . . . . . . . . . . . . . . . . . . 15 blank characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 bn GREP_COLORS capability . . . . . . . . . . . . . . . . . . . . . 11 braces, first argument omitted. . . . . . . . . . . . . . . . . . 14 braces, one argument . . . . . . . . . . . . . . . . . . . . . . . . . . 14 braces, second argument omitted . . . . . . . . . . . . . . . 14 braces, two arguments . . . . . . . . . . . . . . . . . . . . . . . . . 14 bracket expression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Bugs, known . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 bugs, reporting . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 byte offset . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

C

case insensitive search . . . . . . . . . . . . . . . . . . . . . . . 2, 25 changing name of standard input . . . . . . . . . . . . . . . . 5 character class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 character classes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 character encoding . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 19 character type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 classes of characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 cntrl character class. . . . . . . . . . . . . . . . . . . . . . . . . . . 15 context lines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 6 context lines, after match . . . . . . . . . . . . . . . . . . . . . . . 6 context lines, before match . . . . . . . . . . . . . . . . . . . . . . 6 control characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 copying . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 counting lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 cx GREP_COLORS capability . . . . . . . . . . . . . . . . . . . . . 10

D

device search. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 digit character class. . . . . . . . . . . . . . . . . . . . . . . . . . . 15 digit characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 directory search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 dot . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

E

encoding error . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 environment variables . . . . . . . . . . . . . . . . . . . . . . . . . . 10 exclude directories . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 exclude files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 exit status . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

F

FAQ about grep usage . . . . . . . . . . . . . . . . . . . . . . . . . 21 files which don’t match. . . . . . . . . . . . . . . . . . . . . . . . . . 4 fn GREP_COLORS capability . . . . . . . . . . . . . . . . . . . . . 11

G

graph character class. . . . . . . . . . . . . . . . . . . . . . . . . . . 15 graphic characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 grep programs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 GREP_COLOR environment variable. . . . . . . . . . . . . . . 10 GREP_COLORS environment variable . . . . . . . . . . . . . 10 group separator . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

H

hexadecimal digits . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 highlight markers . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 highlight, color, colour . . . . . . . . . . . . . . . . . . . . . . . . . . 4 holes in files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25

I

include files . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 interval expressions. . . . . . . . . . . . . . . . . . . . . . . . . 14, 25 invalid regular expressions. . . . . . . . . . . . . . . . . . . . . . 18 invert matching . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

L

LANG environment variable . . . . . . . . . . . . . . . . . . . 9, 12 language of messages. . . . . . . . . . . . . . . . . . . . . . . . . . . 12 LANGUAGE environment variable . . . . . . . . . . . . . . 9, 12 LC_ALL environment variable. . . . . . . . . . . . . . . . . 9, 12 LC_COLLATE environment variable. . . . . . . . . . . . . . . 12 LC_CTYPE environment variable . . . . . . . . . . . . . . . . . 12 LC_MESSAGES environment variable . . . . . . . . . . . 9, 12 line buffering. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 line numbering . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 ln GREP_COLORS capability . . . . . . . . . . . . . . . . . . . . . 11 locales . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 lower character class. . . . . . . . . . . . . . . . . . . . . . . . . . . 15 lower-case letters. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

Index 39

M

match expression at most m times. . . . . . . . . . . . . . 14 match expression at most once . . . . . . . . . . . . . . . . . 14 match expression from n to m times . . . . . . . . . . . . 14 match expression n or more times . . . . . . . . . . . . . . 14 match expression n times . . . . . . . . . . . . . . . . . . . . . . 14 match expression one or more times . . . . . . . . . . . . 14 match expression zero or more times . . . . . . . . . . . 14 match the whole line . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 matching basic regular expressions . . . . . . . . . . . . . 13 matching extended regular expressions . . . . . . . . . 13 matching fixed strings. . . . . . . . . . . . . . . . . . . . . . . . . . 13 matching Perl-compatible regular expressions … 13 matching whole words. . . . . . . . . . . . . . . . . . . . . . . . . . . 3 max-count . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 mc GREP_COLORS capability . . . . . . . . . . . . . . . . . . . . . 11 message language . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 ms GREP_COLORS capability . . . . . . . . . . . . . . . . . . . . . 11 MS-Windows binary I/O . . . . . . . . . . . . . . . . . . . . . . . . 9 mt GREP_COLORS capability . . . . . . . . . . . . . . . . . . . . . 11

N

names of matching files . . . . . . . . . . . . . . . . . . . . . . . . . 4 national language support. . . . . . . . . . . . . . . . . . . . . . 12 ne GREP_COLORS capability . . . . . . . . . . . . . . . . . . . . . 11 NLS . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 no filename prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 non-ASCII matching . . . . . . . . . . . . . . . . . . . . . . . . . . . 20 non-printable matching . . . . . . . . . . . . . . . . . . . . . . . . 20 null character . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 numeric characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

O

only matching. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 option delimiter . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 ordinary characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

P

patterns from file . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 patterns option . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 performance . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 25 period . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 pipelines and reading . . . . . . . . . . . . . . . . . . . . . . . . . . 25 plus sign . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 POSIXLY_CORRECT environment variable . . . . . . . . . 12 print character class. . . . . . . . . . . . . . . . . . . . . . . . . . . 15 print non-matching lines . . . . . . . . . . . . . . . . . . . . . . . . 3 printable characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 punct character class. . . . . . . . . . . . . . . . . . . . . . . . . . . 15 punctuation characters . . . . . . . . . . . . . . . . . . . . . . . . . 15

Q

question mark . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 quiet, silent . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

R

range expression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 recursive search. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8, 9 regular expressions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 return status. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 rv GREP_COLORS capability . . . . . . . . . . . . . . . . . . . . . 11

S

searching directory trees . . . . . . . . . . . . . . . . . . . . . . 8, 9 searching for patterns . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 sl GREP_COLORS capability . . . . . . . . . . . . . . . . . . . . . 10 space character class. . . . . . . . . . . . . . . . . . . . . . . . . . . 16 space characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 special characters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 subexpression . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 suppress binary data . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 suppress error messages . . . . . . . . . . . . . . . . . . . . . . . . . 5 symbolic links . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8, 9

T

tab-aligned content lines . . . . . . . . . . . . . . . . . . . . . . . . 6 TERM environment variable . . . . . . . . . . . . . . . . . . . . . 12 translation of message language . . . . . . . . . . . . . . . . 12

U

unspecified behavior in regular expressions . . . . . 18 upper character class. . . . . . . . . . . . . . . . . . . . . . . . . . . 16 upper-case letters . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 usage summary, printing . . . . . . . . . . . . . . . . . . . . . . . . 2 usage, examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 using grep, Q&A . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21

V

variants of grep . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 version, printing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

W

whitespace characters . . . . . . . . . . . . . . . . . . . . . . . . . . 16 with filename prefix . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

X

xdigit character class . . . . . . . . . . . . . . . . . . . . . . . . . 16 xdigit class . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

Index 40

Z

zero-terminated file names . . . . . . . . . . . . . . . . . . . . . . 6 zero-terminated lines . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9